OpenStudy (anonymous):

What is the simplified form of y^2 + 2y - 8 over y^2 + 7y + 12

OpenStudy (anonymous):

the top = (y+4)(y-2) and the bottom = (y+4)(y+3). the (y+4)s cancel out soo. you are left with (y-2)/(y+3).